# Eigenvector decorrelation for random matrices

### v2

@@ Abstract.
Giorgio Cipolloni [^1] Department of Mathematics, University of Arizona, 617 N Santa Rita Ave, Tucson, AZ 85721, USA. gcipolloni@arizona.edu László Erdős [^2] Institute of Science and Technology Austria, Am Campus 1, 3400 Klosterneuburg, Austria. lerdos@ist.ac.at Joscha Henheik [^2] Institute of Science and Technology Austria, Am Campus 1, 3400 Klosterneuburg, Austria. joscha.henheik@ist.ac.at Oleksii Kolupaiev [^2] Institute of Science and Technology Austria, Am Campus 1, 3400 Klosterneuburg, Austria. okolupaiev@ist.ac.at 22footnotetext: Supported by the ERC Advanced Grant “RMTBeyond” No. 101020331.
@@ 1.1. Main result
The behavior of the eigenvectors of a Hermitian matrix under perturbations is known to be quite subtle: even a small change in the matrix may lead to a significant rotation of the eigenvectors due to resonances. This phenomenon is ubiquitous in a broad range of numerical and statistical applications, see, e.g., [24, 42, 43, 57, 66, 75]. For example, in the classical paper [42] Davis and Kahan give a deterministic upper bound for the deviation of the eigenvectors from the unperturbed ones in terms of the spectral gap, while in [66, Theorem 2] the authors show that there always exists a perturbation causing a big change in the eigenvectors. When the magnitude of the perturbation exceeds the local eigenvalue spacing of the initial matrix, standard perturbation theory does not control the eigenbasis of the perturbed matrix any more and the behavior of the eigenvectors is highly sensitive to the properties of the original matrix. While in some rare cases even such larger perturbations still cause only a small change, typically the perturbed eigenbasis is completely decoupled from the initial one. In this paper, we show that indeed this typical scenario occurs for random matrices with very high probability.
@@
for bulk indices [^3] We say that an index $i$ is in the bulk of the spectrum if the density of states around $\lambda_{i}^{l}$ is strictly positive; see (2.7) for the precise definition. $i,j$, where $\langle X\rangle:=\frac{1}{N}{\rm Tr}\,X$ denotes the averaged trace of $X\in\mathbf{C}^{N\times N}$. Here $V$ is an appropriately chosen deterministic matrix depending on the deformations $D_{1},D_{2}$ and the (typical locations of the) energies $\lambda_{i}^{1},\lambda_{j}^{2}$ with $\lVert V\rVert\lesssim 1$ (see (2.11) for its definition). The $N^{-1/2}$ error term in (1.1) is optimal.
@@ 1.2. Previous related results
where $\bm{u}_{i}$ are the orthonormal eigenvectors of an $N\times N$ random matrix $H$. The ETH in the form (1.3) was first proven for Wigner matrices (i.e. $D_{1}=D_{2}=0$, in which case $V=I$) in [31] (see also [20, 21] for previous partial results). We point out that even the Gaussianity of the fluctuations in the $N^{-1/2}$-term is known for Wigner matrices for special observables [20, 21], for general observables [33, 12, 11], and for deformed Wigner matrices [27]. The result (1.3) was extended in several directions: to more general random matrix ensembles [27, 1, 74, 52], where $V$ becomes energy dependent, to $d$–regular graphs [10, 9] and to improvement of the error term in (1.3) from $\lVert A\rVert$ to $\langle A^{2}\rangle^{1/2}$ [13, 36, 25]. Related to (1.3), we also mention that in the past few years there has been great interest in studying eigenvector overlaps of different nature in several other contexts, including tensor principal component analysis (PCA) [70], shrinkage estimators [45, 63], noise detection [5, 22], minors [7], the equipartition principle [8], and many body physics [41].
@@ 1.3. Multi–resolvent local laws
Local laws in general are concentration estimates for a single resolvent $G$ of a random matrix, or alternating chains of resolvents and deterministic matrices $A$, i.e. $GAGAGA...$. The main technical tool that we use to prove the decorrelation estimates for eigenvectors in (1.1)–(1.2) is a two-resolvent local law, which is stated in Theorem 3.2 below.
@@
where the $\mathrm{LT}\geq 0$ behaves as (the absolute value of) a linear combination of $D_{1}-D_{2}$ and $z_{1}-z_{2}$ (a precise definition will be given in (2.17) later). The interesting regime is when $\gamma\ll 1$. However, when $A\in\mathbf{C}^{N\times N}$ lies in a specific subspace of codimension one, the bound in (1.5) improves to $\lVert M_{12}^{A}\rVert\lesssim 1$. We call such matrices regular and establish an improved local law for $G_{1}AG_{2}$ in this case. When one deals with Wigner matrices, i.e. $D_{1}=D_{2}=0$, then $A$ is regular if and only if ${\rm Tr}\,A=0$. However, when the deformations $D_{1},D_{2}$ are non–zero, the notion of regularity depends on $D_{1},D_{2}$, as well as on the spectral parameters $z_{1},z_{2}$, in a nontrivial way; see Definition 2.2 for the precise definition.
[…]
@@ 1.4. The method of characteristics
We set $[k]:=\{1,...,k\}$ for $k\in\mathbf{N}$ and $\langle A\rangle:=N^{-1}\mathrm{Tr}(A)$, $N\in\mathbf{N}$, for the normalized trace of an $N\times N$-matrix $A$. For positive quantities $f,g$ we write $f\lesssim g$, $f\gtrsim g$, to denote that $f\leq Cg$ and $f\geq cg$, respectively, for some $N$-independent constants $c,C>0$ that depend only on the basic control parameters of the model in Assumption 2.1 below. We denote the complex upper–half plane by $\mathbf{H}:=\{z\in\mathbf{C}:\Im z>0\}$
@@
Matrix entries are indexed by lower case Roman letters $a,b,c,...,i,j,k,...$ from the beginning or the middle of the alphabet and unrestricted sums over those are always understood to be over $\{1,...,N\}$.
@@ 2. Main results
We start by introducing the notion of regular observables, a concept, which in this generality was first introduced in [28, Def. 3.1] and later in [27, Def. 4.2].
@@
The convenience of our choice of $V$ and thus the definition of regular observables in (2.10) lies in the fact that $V$ is the right eigenvector $R_{12}=R(\nu_{1},\nu_{2})$ corresponding to the smallest (in absolute value) eigenvalue of the operator $\mathcal{X}_{12}$, which is defined by
@@
where $V=V(\nu_{1},\nu_{2})$ is defined in (2.11) and satisfies $\|V\|\lesssim 1$. Here, for a fixed small $\delta=\delta(\kappa)>0$, we defined
(2.15) $\phi_{ij}=\phi_{ij}(\delta):=\bm{1}(|\gamma_{i}^{D_{1}}-\gamma_{j}^{D_{2}}|\leq\delta)\bm{1}(\langle(D_{1}-D_{2})^{2}\rangle\leq\delta).$
The bounds (2.13) and (2.14) are uniform in the indices $i,j$ such that $\gamma_{i}^{D_{1}}\in\mathbf{B}_{\kappa}(D_{1})$ and $\gamma_{j}^{D_{2}}\in\mathbf{B}_{\kappa}(D_{2})$.
@@
    For $D_{1}=D_{2}=0$ (2.14) is the ETH bound for Wigner matrices [31, Theorem 2.2], as in this case $V=I$, yielding
@@
    with $\gamma_{i}:=\gamma_{i}^{D}$. In this case, (2.14) is the ETH bound for deformed Wigner matrices as given in [27, Theorem 2.7]:
@@
Here, $M_{l}=M^{D_{l}}(z_{l})$, for $l=1,2$, is the unique solution [53, Theorem 2.1] of the MDE (1.4) under the constraint $\Im M_{l}\Im z_{l}>0$. We also mention that from (1.4) one can recover (2.5) by $m_{D_{l}}(z_{l})=\langle M_{l}(z_{l})\rangle$. From the definition (2.17) and the fact that $M_{l}$ and $D_{l}$ commute it follows that $\mathrm{LT}(z_{1},z_{2})=\mathrm{LT}(z_{1},\bar{z}_{2})$ and $\mathrm{LT}(z_{1},z_{2})=\mathrm{LT}(\bar{z}_{1},z_{2})$ for any $z_{1},z_{2}\in\mathbf{C}\setminus\mathbf{R}$. Therefore, (2.17) extends continuously to the real line, i.e. $\mathrm{LT}(z_{1},z_{2})$ is well-defined for $z_{1},z_{2}\in\mathbf{R}$.
@@
$\mathbf{P}^{\rm GOE/GUE}\big{(}|\lambda_{i}^{1}-\lambda_{i+1}^{1}|\leq N^{-1-\omega}\big{)}\leq N^{-c\omega}.$
@@
We point out that using the eigenvector overlap bound (2.18) we can prove that the eigenvalue gaps in the bulk of the spectrum of $W+D_{1}$, $W+D_{2}$ are independent as long as $\langle(D_{1}-D_{2})^{2}\rangle\gg 1/N$. In fact, following verbatim [29, Section 7] and its adaptation to the Hermitian case in [35], we can prove the desired independence via the study of weakly correlated Dyson Brownian motions. The only input required for this proof is the overlap bound $|\langle{\bm{u}}_{i}^{1},{\bm{u}}_{j}^{2}\rangle|\ll 1$.
@@ 3. Proofs of the Main Results: Multi-resolvent local laws
In this section we present several technical tools and preliminary results that will be often used in this paper. More precisely, in Section 3.1 we study lower bounds on the stability operator, which are one of the fundamental input to obtain the decay in the rhs. of (2.18). Then, in Section 3.2, we state our main technical result (Theorem 3.2 below), which is a multi–resolvent local law for the product of the resolvents of $W+D_{1}$ and $W+D_{2}$, with $D_{1},D_{2}\in\mathbf{C}^{N\times N}$. Lastly, in Section 3.3 we comment on the optimality and discuss some possible extension of Theorem 3.2.
@@ 3.1. Preliminaries on the stability operator
The main control parameters in the following statements are $\langle(D_{1}-D_{2})^{2}\rangle$ and the linear term $\mathrm{LT}(z_{1},z_{2})$ which is defined as in (2.17), for $z_{1},z_{2}\in\mathbf{C}\setminus\mathbf{R}$. The proof of the following proposition and comments about its optimality are postponed to Section A.1.
@@
The most relevant part of Proposition 3.1 is the lower bound $\beta_{*}\gtrsim\widehat{\gamma}$. This bound in a weaker form (more precisely, without $\mathrm{LT}$ included in $\widehat{\gamma}$) has already appeared in [48, Proposition 4.2]. It should be viewed as an upper bound on the two–body stability operator in terms of simpler control parameters collected in $\widehat{\gamma}$. In the inequality $\beta_{*}\lesssim\widehat{\gamma}^{1/4}$ we do not pursue getting the optimal power for $\widehat{\gamma}$. In fact, any positive exponent would work for our purpose.
@@
Let $D\in\mathbf{C}^{N\times N}$, with $\lVert D\rVert\lesssim 1$, and let $W$ be a Wigner matrix satisfying Assumption 2.1. Then, for $z\in\mathbf{C}\setminus\mathbf{R}$ we define the resolvent of $W+D$ by $G(z)=G^{D}(z):=(W+D-z)^{-1}$. It is well known that in the limit $N\to\infty$ the resolvent becomes approximately deterministic $G(z)\approx M(z)$, with $M(z)=M^{D}(z)$ being the solution of (1.4). This is expressed by the following single resolvent local law [51, Theorem 2.1]
@@
We first consider the bound (3.11). In the bulk of the spectrum, this bound is optimal except for the fact that $1/(\sqrt{N\eta_{*}}\widehat{\gamma})$ should be replaced by $1/(N\eta_{*}\widehat{\gamma})$. Notice, that once the bound $1/(N\eta_{*}\widehat{\gamma})$ is achieved, the term $1/(N\eta_{1}\eta_{2})$ in (3.11) is obsolete, as it is always bigger. This improvement can be achieved by proving (weaker) local laws also for products of longer resolvents; see, e.g., [34, 25, 26] for similar arguments. In fact, this overestimate is due to the fact that the four resolvents chains, appearing e.g. in the quadratic variation of the stochastic term in (5.2) below are currently estimated in terms of products of traces of two resolvents using certain crude reduction inequalities (see e.g. (5.27)). Following the evolution of these longer chains more carefully would give the improvement $1/(N\eta_{*}\widehat{\gamma})$.
@@ 4.1. Input global laws
The proof of these global laws is very similar to the one presented in [34, Appendix B], we thus omit several details and just present the main steps. To keep the presentation short and simple we only present the proof in the averaged case.
@@
Finally, using a minimalistic cumulant expansion as in [34, (B.4)–(B.8)], we conclude $|\langle M_{1}\underline{WG_{1}B_{1}G_{2}B_{2}}\rangle|\prec N^{-1}\lVert B_{1}\rVert\lVert B_{2}\rVert$ and so (4.2). ∎

### v3

@@ Abstract.
Giorgio Cipolloni [^1] Department of Mathematics, University of Rome Tor Vergata, Via della Ricerca Scientifica, 1, 00133 Roma RM, Italy. cipolloni@axp.mat.uniroma2.it László Erdős [^2] Institute of Science and Technology Austria, Am Campus 1, 3400 Klosterneuburg, Austria. lerdos@ist.ac.at Joscha Henheik [^3] Section of Mathematics, University of Geneva, Rue du Conseil Général 7-9, 1205 Geneva, Switzerland. joscha.henheik@unige.ch Oleksii Kolupaiev [^2] Institute of Science and Technology Austria, Am Campus 1, 3400 Klosterneuburg, Austria. okolupaiev@ist.ac.at
@@ 1.1. Main result
The behavior of the eigenvectors of a Hermitian matrix under perturbations is known to be quite subtle: even a small change in the matrix may lead to a significant rotation of the eigenvectors due to resonances. This phenomenon is ubiquitous in a broad range of numerical and statistical applications, see, e.g., [24, 43, 44, 58, 67, 76]. For example, in the classical paper [43] Davis and Kahan give a deterministic upper bound for the deviation of the eigenvectors from the unperturbed ones in terms of the spectral gap, while in [67, Theorem 2] the authors show that there always exists a perturbation causing a big change in the eigenvectors. When the magnitude of the perturbation exceeds the local eigenvalue spacing of the initial matrix, standard perturbation theory does not control the eigenbasis of the perturbed matrix any more and the behavior of the eigenvectors is highly sensitive to the properties of the original matrix. While in some rare cases even such larger perturbations still cause only a small change, typically the perturbed eigenbasis is completely decoupled from the initial one. In this paper, we show that indeed this typical scenario occurs for random matrices with very high probability.
@@
for bulk indices [^3] We say that an index $i$ is in the bulk of the spectrum if the density of states around $\lambda_{i}^{l}$ is strictly positive; see (2.7) for the precise definition. $i,j$, where $\langle X\rangle:=\frac{1}{N}{\mathrm{T}r}\,X$ denotes the averaged trace of $X\in\mathbf{C}^{N\times N}$. Here $V$ is an appropriately chosen deterministic matrix depending on the deformations $D_{1},D_{2}$ and the (typical locations of the) energies $\lambda_{i}^{1},\lambda_{j}^{2}$ with $\lVert V\rVert\lesssim 1$ (see (2.11) for its definition). The $N^{-1/2}$ error term in (1.1) is optimal.
@@ 1.2. Previous related results
where $\bm{u}_{i}$ are the orthonormal eigenvectors of an $N\times N$ random matrix $H$. The ETH in the form (1.3) was first proven for Wigner matrices (i.e. $D_{1}=D_{2}=0$, in which case $V=I$) in [32] (see also [20, 21] for previous partial results). We point out that even the Gaussianity of the fluctuations in the $N^{-1/2}$-term is known for Wigner matrices for special observables [20, 21], for general observables [34, 12, 11], and for deformed Wigner matrices [28]. The result (1.3) was extended in several directions: to more general random matrix ensembles [28, 1, 75, 53], where $V$ becomes energy dependent, to dd–regular graphs [10, 9] and to improvement of the error term in (1.3) from $\lVert A\rVert$ to $\langle A^{2}\rangle^{1/2}$ [13, 37, 25]. Related to (1.3), we also mention that in the past few years there has been great interest in studying eigenvector overlaps of different nature in several other contexts, including tensor principal component analysis (PCA) [71], shrinkage estimators [46, 64], noise detection [5, 22], minors [7], the equipartition principle [8], and many body physics [42].
@@ 1.3. Multi–resolvent local laws
Local laws in general are concentration estimates for a single resolvent $G$ of a random matrix, or alternating chains of resolvents and deterministic matrices $A$, i.e. GAGAGAGAGAGA.... The main technical tool that we use to prove the decorrelation estimates for eigenvectors in (1.1)–(1.2) is a two-resolvent local law, which is stated in Theorem 3.2 below.
@@
where the $\mathrm{LT}\geq 0$ behaves as (the absolute value of) a linear combination of $D_{1}-D_{2}$ and $z_{1}-z_{2}$ (a precise definition will be given in (2.17) later). The interesting regime is when $\gamma\ll 1$. However, when $A\in\mathbf{C}^{N\times N}$ lies in a specific subspace of codimension one, the bound in (1.5) improves to $\lVert M_{12}^{A}\rVert\lesssim 1$. We call such matrices regular and establish an improved local law for $G_{1}AG_{2}$ in this case. When one deals with Wigner matrices, i.e. $D_{1}=D_{2}=0$, then $A$ is regular if and only if ${\mathrm{T}r}\,A=0$. However, when the deformations $D_{1},D_{2}$ are non–zero, the notion of regularity depends on $D_{1},D_{2}$, as well as on the spectral parameters $z_{1},z_{2}$, in a nontrivial way; see Definition 2.2 for the precise definition.
We now informally discuss the structure of the bounds in the multi–resolvent local laws in Theorem 3.2 and Proposition 4.16 with a concrete example. Let $\bm{x},\bm{y}\in\mathbf{C}^{N}$ be deterministic unit vectors. When $D_{1}=D_{2}$, it was shown in [28, Proposition 4.4] that for $\lVert A\rVert\lesssim 1$ we have [^5] By $\langle\cdot,\cdot\rangle$ we denote the inner product in $\mathbf{C}^{N}$.
@@ 1.4. The method of characteristics
We set $k]:=\{1,...,k\}$ for $k\in\mathbf{N}$ and $\langle A\rangle:=N^{-1}\mathrm{Tr}(A)$, $N\in\mathbf{N}$, for the normalized trace of an $N\times N$-matrix $A$. For positive quantities $f,g$ we write $f\lesssim g$, $f\gtrsim g$, to denote that $f\leq Cg$ and $f\geq cg$, respectively, for some $N$-independent constants $c,C>0$ that depend only on the basic control parameters of the model in Assumption 2.1 below. We denote the complex upper–half plane by $\mathbf{H}:=\{z\in\mathbf{C}:\Im z>0\}$
@@
Matrix entries are indexed by lower case Roman letters a,b,c,,i,j,k,a,b,c,...,i,j,k,... from the beginning or the middle of the alphabet and unrestricted sums over those are always understood to be over $\{1,...,N\}$.
@@
Acknowledgments
L. Erdős, J. Henheik and O. Kolupaiev were supported by the ERC Advanced Grant “RMTBeyond” No. 101020331. G. Cipolloni is partially supported by the MUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C18000100006.
@@ 2. Main results
We start by introducing the notion of regular observables, a concept, which in this generality was first introduced in [29, Def. 3.1] and later in [28, Def. 4.2].
@@
The convenience of our choice of $V$ and thus the definition of regular observables in (2.10) lies in the fact that $V$ is the right eigenvector $R_{12}=R(\nu_{1},\nu_{2})$ corresponding to the largest (in absolute value) eigenvalue of the operator $\mathcal{X}_{12}$, which is defined by
@@
where $V=V(\nu_{1},\nu_{2})$ is defined in (2.11). Here we defined
(2.15) $\phi_{ij}=\phi_{ij}(\delta):=\bm{1}(|\gamma_{i}^{D_{1}}-\gamma_{j}^{D_{2}}|\leq\delta)\bm{1}(\langle(D_{1}-D_{2})^{2}\rangle\leq\delta)$
for a fixed $\delta=\delta(\kappa)>0$ which is chosen sufficiently small, so that $\|V\|\lesssim 1$ when $\phi_{ij}\neq 0$. The bounds (2.13) and (2.14) are uniform in the indices $i,j$ such that $\gamma_{i}^{D_{1}}\in\mathbf{B}_{\kappa}(D_{1})$ and $\gamma_{j}^{D_{2}}\in\mathbf{B}_{\kappa}(D_{2})$.
@@
    For $D_{1}=D_{2}=0$ (2.14) is the ETH bound for Wigner matrices [32, Theorem 2.2], as in this case $V=I$, yielding
@@
    with $\gamma_{i}:=\gamma_{i}^{D}$. In this case, (2.14) is the ETH bound for deformed Wigner matrices as given in [28, Theorem 2.7]:
@@
Here, $M_{l}=M^{D_{l}}(z_{l})$, for $l=1,2$, is the unique solution [54, Theorem 2.1] of the MDE (1.4) under the constraint $\Im M_{l}\Im z_{l}>0$. We also mention that from (1.4) one can recover (2.5) by $m_{D_{l}}(z_{l})=\langle M_{l}(z_{l})\rangle$. From the definition (2.17) and the fact that $M_{l}$ and $D_{l}$ commute it follows that $\mathrm{LT}(z_{1},z_{2})=\mathrm{LT}(z_{1},\bar{z}_{2})$ and $\mathrm{LT}(z_{1},z_{2})=\mathrm{LT}(\bar{z}_{1},z_{2})$ for any $z_{1},z_{2}\in\mathbf{C}\setminus\mathbf{R}$. Therefore, (2.17) extends continuously to the real line, i.e. $\mathrm{LT}(z_{1},z_{2})$ is well-defined for $z_{1},z_{2}\in\mathbf{R}$.
@@
$\mathbf{P}^{\mathrm{G}OE/GUE}\big(|\lambda_{i}^{1}-\lambda_{i+1}^{1}|\leq N^{-1-\omega}\big)\leq N^{-c\omega}.$
@@
We point out that using the eigenvector overlap bound (2.18) we can prove that the eigenvalue gaps in the bulk of the spectrum of $W+D_{1}$, $W+D_{2}$ are independent as long as $\langle(D_{1}-D_{2})^{2}\rangle\gg 1/N$. In fact, following verbatim [30, Section 7] and its adaptation to the Hermitian case in [36], we can prove the desired independence via the study of weakly correlated Dyson Brownian motions. The only input required for this proof is the overlap bound $|\langle{\bm{u}}_{i}^{1},{\bm{u}}_{j}^{2}\rangle|\ll 1$.
@@ 3. Proofs of the Main Results: Multi-resolvent local laws
In this section we present several technical tools and preliminary results that will be often used in this paper. More precisely, in Section 3.1 we study lower bounds on the stability operator, which are one of the fundamental inputs to obtain the decay in the rhs. of (2.18). Then, in Section 3.2, we state our main technical result (Theorem 3.2 below), which is a multi–resolvent local law for the product of the resolvents of $W+D_{1}$ and $W+D_{2}$, with $D_{1},D_{2}\in\mathbf{C}^{N\times N}$. Lastly, in Section 3.3 we comment on the optimality and discuss some possible extension of Theorem 3.2.
@@ 3.1. Preliminaries on the stability operator
The main control parameters in the following statements are $\langle(D_{1}-D_{2})^{2}\rangle$ and the linear term $\mathrm{LT}(z_{1},z_{2})$ which is defined as in (2.17), for $z_{1},z_{2}\in\mathbf{C}\setminus\mathbf{R}$. The proof of the following proposition and comments about its optimality are postponed to Section A.1 of the Supplementary Material [27].
@@
The most relevant part of Proposition 3.1 is the lower bound $\beta_{*}\gtrsim\widehat{\gamma}$. This bound in a weaker form (more precisely, without $\mathrm{LT}$ included in $\widehat{\gamma}$) has already appeared in [49, Proposition 4.2]. It should be viewed as a lower bound on the two–body stability operator in terms of simpler control parameters collected in $\widehat{\gamma}$. In the inequality $\beta_{*}\lesssim\widehat{\gamma}^{1/4}$ we do not pursue getting the optimal power for $\widehat{\gamma}$. In fact, any positive exponent would work for our purpose.
@@
Let $D\in\mathbf{C}^{N\times N}$, with $\lVert D\rVert\lesssim 1$, and let $W$ be a Wigner matrix satisfying Assumption 2.1. Then, for $z\in\mathbf{C}\setminus\mathbf{R}$ we define the resolvent of $W+D$ by $G(z)=G^{D}(z):=(W+D-z)^{-1}$. It is well known that in the limit $N\to\infty$ the resolvent becomes approximately deterministic $G(z)\approx M(z)$, with $M(z)=M^{D}(z)$ being the solution of (1.4). This is expressed by the following single resolvent local law [52, Theorem 2.1]
@@
[…]
@@ 4.1. Input global laws
The proof of these global laws is very similar to the one presented in [35, Appendix B], we thus omit several details and just present the main steps. To keep the presentation short and simple we only present the proof in the averaged case.
@@
Finally, using a minimalistic cumulant expansion as in [35, (B.4)–(B.8)], we conclude $|\langle M_{1}\underline{WG_{1}B_{1}G_{2}B_{2}}\rangle|\prec N^{-1}\lVert B_{1}\rVert\lVert B_{2}\rVert$ and so (4.2). ∎